OpenStudy (anonymous):

[T] 9. Write the equation of a parabola in factored form that has a step pattern of -3, -9, -15… and has zeroes of 2 and -5.

OpenStudy (anonymous):

step pattern?

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

zeros of 2 and -5 means you can start with \[y=a(x-2)(x+5)=a(x^2+3x-10)\]

OpenStudy (anonymous):

how did you get a(x2+3x−10)?

OpenStudy (anonymous):

because if you know the zeros, you know how to factor

OpenStudy (anonymous):

im sorry but i dont understand ! that was the question straight from my homework !

OpenStudy (anonymous):

just like if i was asked to solve \[x^2+3x-10=0\] i would factor as \((x+5)(x-2)=0\) and get \(x=-5,x=2\) now we go the other way

OpenStudy (anonymous):

you know the zeros, so you know how to factor the zeros are \(-5, 2\) so it must factor as \(a(x+5)(x-2)=a(x^2+3x-10)\)

OpenStudy (anonymous):

damned if i know what a step patter is though do you?

OpenStudy (anonymous):

OHHH so you expand!